OpenStudy (anonymous):

Multiply: (x - 4)(x - 5).

OpenStudy (anonymous):

four multiplications, then combine like terms

OpenStudy (anonymous):

\[(x-4)(x-5)=x\times x-4\times x-5\times x+(-5)\times (-4)\] is a start

OpenStudy (anonymous):

then \[x^2-4x-5x+20\] and finally \[x^2-9x+20\]